An electrochemical microsensor can detect single nanoparticles colliding on its surface. The average collision frequency measured in an experiment is 1.3 particles/s. What is the estimated probably of observing two separate particles on the same sensor in 0.1 second?
Answer #1

expected number of particles in 0.1 seconds =1.3*0.1 =0.13

therefore from Poisson distribution:

estimated probability of observing two separate particles on the same sensor in 0.1 second

=P(X=2)=(e-0.13*0.132/2!) =0.0074